Cane Company manufactures two products called Alpha and Beta that sell for $170 and $130, respectively. Each product uses only one type of raw material that costs $6 per pound. The company has the capacity to annually produce 116,000 units of each product. Its average cost per unit for each product at this level of activity are given below:

Alpha Beta
Direct materials $ 30 $ 18
Direct labor 30 25
Variable manufacturing overhead 20 15
Traceable fixed manufacturing overhead 26 28
Variable selling expenses 22 18
Common fixed expenses 25 20
Total cost per unit $ 153 $ 124

The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are unavoidable and have been allocated to products based on sales dollars.

Required:

1. What is the total amount of traceable fixed manufacturing overhead for each of the two products?

2. What is the company’s total amount of common fixed expenses?

3. Assume that Cane expects to produce and sell 90,000 Alphas during the current year. One of Cane's sales representatives has found a new customer who is willing to buy 20,000 additional Alphas for a price of $120 per unit. What is the financial advantage (disadvantage) of accepting the new customer's order?

4. Assume that Cane expects to produce and sell 100,000 Betas during the current year. One of Cane’s sales representatives has found a new customer who is willing to buy 3,000 additional Betas for a price of $49 per unit. What is the financial advantage (disadvantage) of accepting the new customer's order?

5. Assume that Cane expects to produce and sell 105,000 Alphas during the current year. One of Cane's sales representatives has found a new customer who is willing to buy 20,000 additional Alphas for a price of $120 per unit; however pursuing this opportunity will decrease Alpha sales to regular customers by 9,000 units.

a. What is the financial advantage (disadvantage) of accepting the new customer’s order?

b. Based on your calculations above should the special order be accepted?
Answer #1

Solution 1:

Total amount of traceable fixed manufacturing overhead for the Alpha product line = 116000*$26 = $3,016,000

Total amount of traceable fixed manufacturing overhead for the Beta product line = 116000*$28 = $3,248,000

Solution 2:

Total amount of common fixed expenses = ($25 + $20) * 116000 = $5,220,000

Solution 3:

Contribution margin per unit on special order of alpha= Selling price - Variable cost per unit

= $120 - ($30 + $30 + $20 + $12) = $28 per unit

If Cane accepts the customer’s offer for 200000 alpha, then increase in operating income = 20000*$28 = $560,000

Solution 4:

Contribution margin per unit on special order of Beta= Selling price - Variable cost per unit

= $49 - ($18 + $25 + $15 + $18) = -$27

If Cane accepts the customer’s offer for 3000 Beta, then increase (decrease) in operating income = 3000*(-$27) = -$81,000

Solution 5:

Regular contribution margin per unit for Alpha = $170 - ($30 + $30 + $20 + $22) = $78 per unit

Contribution margin per unit on special order of alpha= Selling price - Variable cost per unit

= $120 - ($30 + $30 + $20 + $12) = $28 per unit

If Cane accepts the customer’s offer for 20000 alpha, additional contribution margin from special order = 20000*$28 = $560,000

If Cane accept special order then loss of contribution margin on regular order = 9000*$78 = $702,000

Incremental net operating income if the order is accepted = $560,000 - $702,000 = ($142,000)

As there is net financial disadvantage, therefore special order should not be accepted.